1. A carton of 24 light bulbs includes two that are defecetive. If two of the bulbs are chosen at random, what are the probabilities that

a Neither bulb will be defective

b One of the bulbs will be defective

c Both bulbs will be defective



1
Expert's answer
2021-05-18T13:09:33-0400

a)

"P=\frac{22}{24}\cdot\frac{21}{23}=\frac{77}{92}"


b)

"P=\frac{22}{24}\cdot\frac{2}{23}=\frac{11}{138}"


c)

"P=\frac{2}{24}\cdot\frac{1}{23}=\frac{1}{276}"